OpenStudy (anonymous):

Find the derivative. h(z)=e^-z^2/z

OpenStudy (anonymous):

You mean this?\[h(z)=\frac{e ^{-z^2} }{ z}\]

OpenStudy (anonymous):

h(z)=e^(-z^2/z)

OpenStudy (anonymous):

apply the chain rule: \[e ^{\frac{ -z^2 }{ z }}=(\frac{ -z^2 }{ z})'*e ^{\frac{ -z^2 }{ z }}\]

OpenStudy (anonymous):

would you mind showing me how to do this?

OpenStudy (anonymous):

i just did ask me what's unclear

OpenStudy (anonymous):

what would the answer be for this problem? I have two answers but I do not think either one is right

OpenStudy (anonymous):

I will continue even if the answer is almost obvious\[(\frac{ -z^2 }{ z })'*e ^{\frac{ -z^2 }{ z }}=(-z)'*e ^{-z}\]

OpenStudy (anonymous):

okay, thanks that actually was one of my answers